in direct variation if y varies directly with x:
find the equation for the direct variation if y=20 when x= 3 find x when y=8

Answer by drk(1908)

Step 1 write y = kx and find k using known x and y information.
y = kx
20 = 3k
k = 20/3
step 2 - rewrite the equation with your new K value as
y = (20/3)x
step 3 find x when y = 8
8 = (20/3)x
24/20 = x
x = 6/5